OpenStudy (dimples97):

Can someone help me simplify this expression? √169√169

TheSmartOne (thesmartone):

HInt: \(\sqrt{x} \times \sqrt{x} = (\sqrt{x})^2 = x\) So what is \(\sqrt{169} \times \sqrt{169} = (\sqrt{169})^2 = ??\)

OpenStudy (aaronandyson):

Do you still need help?

TheSmartOne (thesmartone):